# iMTSP: Solving Min-Max Multiple Traveling Salesman Problem with Imperative Learning

### v2

@@ I Introduction
To enhance generalization and accelerate the training process, we reformulate the Min-Max MTSP as a bilevel optimization problem. This comprises an upper-level optimization that focuses on training a city allocation network, and a lower-level optimization that solves multiple TSPs, inspired by imperative learning (IL) [24, 25, 26]. Specifically, the allocation network is to assign each city to an agent, which decomposes the MTSP into multiple single-agent TSPs. Then, a classic TSP solver, which can quickly produce a near-optimal tour for a single-agent TSP with up to a few hundred cities, is employed to find a tour for each agent based on the allocation. A key aspect is that the upper-level optimization, namely the training of the allocation network, is supervised by the lower-level TSP solver. This metric-based supervision from the lower-level optimization results in an end-to-end, self-supervised framework for solving MTSP, which we refer to as imperative MTSP (iMTSP).
One technical obstacle when developing iMTSP is that the space of possible city allocations is discrete. As a result, the objective function, i.e., the maximum tour length given by the lower-level TSP solver, is non-differentiable to the city allocation, which prevents the back-propagation to the allocation network’s parameters. One of the solutions is to explore the policy gradient algorithm or the reparameterization trick [27], which often use Monte-Carlo methods to estimate the gradient. However, these methods can lead to gradient estimations with high variance since the space of city allocation is extremely large. To handle this difficulty, we introduce a control variate [28] to reduce the variance of the gradient estimations. Parameterized by a trainable surrogate network running in parallel with the TSP solver, the control variate can be efficiently computed, and the overall algorithm can provide low variance gradient information through the non-differentiable solver and the discrete decision space. The main contributions of this paper are summarized as follows.
@@ II Related works
Heuristic approaches for MTSP handle the challenge by decomposing the original problem into several phases which in turn reduces the computational burden of each phase. Examples are particle swarm optimization [31], ant colony optimization [12], etc [32, 33].
Although heuristic algorithms usually run faster than analytical approaches, they lack theoretic guarantees on solution quality and could produce highly sub-optimal solutions especially for large-scale problems with hundreds and thousands of cities.
@@ III Problem Formulation
Given a set of $N$ cities, each represented by a 2-D coordinate $\mathbf{x_{i}}\in\mathbb{R}^{2}$ and collectively denoted as $X=\{\mathbf{x_{1}},...,\mathbf{x_{N}}\}$, along with $M$ agents starting from a common city, i.e., the depot $\mathbf{x_{d}}\in X$, the MTSP is characterized by the tuple $(X,\mathbf{x_{d}},M)$. The objective is to find, for each agent $j$, a tour $T_{j}=\{\mathbf{x}_{j,1},...,\mathbf{x}_{j,k},...\}$ such that all cities are visited exactly once, each tour returns to the depot, and the length of the longest tour reaches the minimum. The overall optimization objective for Min-Max MTSP can be defined as:
@@
where the tour length $\mathbf{D}$ is the distance between two cities:
@@ IV-A Reformulation of MTSP
where $a_{j}$ is the allocation for agent $j$ sampling the probability matrix by row, and $L$ returns the maximum route length among all agents. $s$ is a surrogate network with parameter $\gamma$ , which will be explained in detail in Section IV-C along with the specific definitions of upper-level objectives. Specifically, the upper-level optimization consists of two separate steps: the upper-level cost $U_{\theta}$ is to optimize the allocation network $f(\theta)$, while $U_{\gamma}$ is to optimize the surrogate network $s(\gamma)$. The framework of iMTSP is shown in Fig. 1.
@@ IV-B Forward Propagation Details
We next present the details of solving iMTSP (3). For each instance with coordinates $X$, we first embed agents and cities into feature vectors that contain the agent-city and city-city topological information. The embedded instance $X^{\prime}$ is the input to the allocation network $P=f(X^{\prime},\mathbf{x_{d}},M;\mathbf{\theta})$, where $\mathbf{\theta}$ is the network parameters. Each element $p_{i,j}$ in the output $P\in\mathbb{R}^{N\times M}$ is the probability that the $i$-th city is allocated to the $j$-th agent. This probability matrix is then sampled in each row so that each city is assigned to an agent. The sampling result is then reorganized into allocations $A=\{a_{1},a_{2},...,a_{M}\}$, where $a_{j}$ represents the set of cities assigned to the $j$-th agent. Subsequently, a TSP solver $g(a_{j};\mathbf{\mu})$ for each agent $j$ is invoked to order the cities in the allocation $a_{j}$ and return a tour. Finally, the cost of each tour is computed with (2) and the length of the longest tour is the cost of that instance, i.e., $L=\max_{j}D(g(a_{j};\mu))$. The internal process of embedding and the allocation network is below.
@@ IV-C2 Control Variate
$h(d)_{\text{new}}=h(d)+\gamma(c(d)-\mathbb{E}[c(d)]).$ (10)
When constant $\gamma$ is properly chosen, $h(d)_{new}$ has the same expected value but lower variance than $h(d)$, as long as the control variate $c(d)$ is correlated to $h(d)$. We briefly summarize the underlying principle of control variate technique [45] as follows and omit the dependency on $d$ for simplicity.
@@
Let $h$ denotes an RV with an unknown expected value $\mathbb{E}(h)$, and $c$ denotes an RV with a known expected value $\mathbb{E}(c)$. The new RV $h_{new}=h+\gamma(c-\mathbb{E}(c))$ has same expected value but smaller variance, when the constant $\gamma$ is properly chosen and $c$ is correlated with $h$.
@@
$\mathrm{Var}(h_{\text{new}})=\mathrm{Var}(h)+\gamma^{2}\mathrm{Var}(c)+2\gamma\mathrm{Cov}(h,c).$ (11)
Differentiate the above equation w.r.t. $\gamma$ and set the derivative to zero, the optimal choice of $\gamma$ is
$\gamma^{*}=-\frac{\mathrm{Cov}(h,c)}{\mathrm{Var}(c)}.$ (12)
The minimum variance of $h_{\text{new}}$ with $\gamma^{*}$ is $(1-\rho^{2}_{h,c})\mathrm{Var}(h)$ where $\rho$ is the correlation coefficient of $h$ and $c$. Since $\rho\in[-1,1]$, the inequality $\mathrm{Var}(h_{\text{new}})\leq\mathrm{Var}(h)$ always holds with $\gamma=\gamma^{*}$. When $\lVert\rho\rVert\rightarrow 1$, $\mathrm{Var}(h_{\text{new}})\rightarrow 0$. ∎
@@
In practice, $\gamma^{*}$ is usually inaccessible, and thus we cannot arbitrarily choose $c$. Without losing generality, consider the case where $h$ and $c$ are positively correlated. By solving the inequality $\gamma^{2}\mathrm{Var}(c)+2\gamma\mathrm{Cov}(h,c)<0$, we get $\gamma\in(\frac{-2\mathrm{Cov}(h,c)}{\mathrm{Var}(c)},0)$, which indicates that as long as $k$ is set in this range, $\mathrm{Var}(h_{\text{new}})$ is less than $\mathrm{Var}(h)$. When $h$ and $c$ are highly correlated and are of similar ranges, $\mathrm{Cov}(h,c)\approx\mathrm{Var}(c)$, thus $\gamma^{*}\approx-1$, which is our empirical choice of $\gamma$.
Inspired by the control variant gradient estimator (10) and to have a lower gradient variance, we design the upper-level costs $U_{\theta}$ for the allocation network as:
@@
where $L^{\prime}$ is output of the surrogate network (control variate) but is taken as a constant by cutting its gradients. In turn, our gradient for the allocation network parameters $\theta$ is:
@@ V Experiments
We next conduct comprehensive experiments to demonstrate the generalization ability, computational efficiency, and convergence speed of our iMTSP framework.
@@ V-A Baselines
To demonstrate the advantages of our method, we compare it with both classic methods and learning-based methods. Specifically, we select a state-of-the-art reinforcement learning (RL)-based model from [14] as a baseline. Their framework can efficiently solve MTSPs with up to a few thousand cities. Using S-batched REINFORCE, it has shown outstanding generalization ability and has a reduced variance for the gradient estimator than the vanilla REINFORCE [23]. The other potential baseline methods [13, 46, 20] either don’t provide source codes or cannot be directly applied to our settings. We also compare iMTSP with a well-known meta-heuristic method implemented by the Google OR-Tools routing library [18, 22, 14]. It provides near-optimal solutions to MTSPs with a few hundred cities and is a popular MTSP baseline [29]. For abbreviation, we will refer to the two methods as an RL baseline and Google OR-Tools.

### v3

@@ I Introduction
To enhance generalization and accelerate the training process, we reformulate the Min-Max MTSP as a bilevel optimization problem. This comprises an upper-level optimization that focuses on training a city allocation network, and a lower-level optimization that solves multiple signal-agent TSPs, inspired by imperative learning (IL) [24, 25, 26]. Specifically, the allocation network is to assign each city to an agent, which decomposes the MTSP into multiple single-agent TSPs. Then, a classic TSP solver, which can quickly produce a near-optimal tour for a single-agent TSP with up to a few hundred cities, is employed to find a tour for each agent based on the allocation. A key aspect is that the upper-level optimization, namely the training of the allocation network, is supervised by the lower-level TSP solver. This metric-based supervision from the lower-level optimization results in an end-to-end, self-supervised framework for solving MTSP, which we refer to as imperative MTSP (iMTSP).
One technical obstacle when developing iMTSP is that the space of possible city allocations is discrete. As a result, the objective function, i.e., the maximum tour length given by the lower-level TSP solver, is non-differentiable to the city allocation, which prevents the back-propagation to the allocation network’s parameters. One of the solutions is to explore the policy gradient algorithm or the reparameterization trick [27], which often use Monte-Carlo methods to estimate the gradient. However, these methods can lead to gradient estimations with high variance since the space of city allocation is extremely large. To handle this difficulty, we introduce a control variate [28] to reduce the variance of the gradient estimations. Parameterized by a trainable surrogate network running in parallel with the TSP solver, the control variate can be efficiently computed, and the overall algorithm can provide low variance gradient information through the non-differentiable solver and the discrete decision space. The main contributions of this paper are summarized as follows:
@@ II Related works
Heuristic approaches for MTSP handle the challenge by decomposing the original problem into several phases which in turn reduces the computational burden of each phase. Examples are particle swarm optimization [31], ant colony optimization [12], etc [32, 33]. Although heuristic algorithms usually run faster than analytical approaches, they lack theoretic guarantees on solution quality and could produce highly sub-optimal solutions especially for large-scale problems with hundreds and thousands of cities.
@@ III Problem Formulation
Given a set of $N$ cities collectively denoted as $X=\{\mathbf{x_{1}},...,\mathbf{x_{N}}\}$, along with $M$ agents starting from a common city, i.e., the depot $\mathbf{x_{d}}\in X$, the MTSP is characterized by the tuple $(X,\mathbf{x_{d}},M)$. The objective is to find, for each agent $j$, a tour $T_{j}=\{\mathbf{x}_{j,1},...,\mathbf{x}_{j,k},...\}$ such that all cities are visited exactly once, each tour returns to the depot, and the length of the longest tour reaches the minimum. The optimization objective for Min-Max MTSP can be defined as:
@@
where $D(T_{j})$ is the cost of route $T_{j}$. For easier understanding, we use Euclidean distance as the cost in this paper, i.e.,:
@@ IV-A Reformulation of MTSP
where $a_{j}$ is the allocation for agent $j$ sampling from the probability matrix by row, and $L$ returns the maximum route length among all agents. $s$ is a surrogate network with parameter $\gamma$ , which will be explained in detail in Section IV-C along with the specific definitions of upper-level objectives. Specifically, the upper-level optimization consists of two separate steps: the upper-level cost $U_{\theta}$ is to optimize the allocation network $f(\theta)$, while $U_{\gamma}$ is to optimize the surrogate network $s(\gamma)$. The framework of iMTSP is shown in Fig. 1.
@@ IV-B Forward Propagation Details
We next present the details of solving iMTSP (3). For each instance with coordinates $X$, we first embed agents and cities into feature vectors that contain the agent-city and city-city topological information. The embedded instance $X^{\prime}$ is the input to the allocation network $P=f(X^{\prime},\mathbf{x_{d}},M;\mathbf{\theta})$. Each element $p_{i,j}$ in the output $P\in\mathbb{R}^{N\times M}$ is the probability that the $i$-th city is allocated to the $j$-th agent. This probability matrix is then sampled in each row so that each city is assigned to an agent. The sampling result is then reorganized into allocations $A=\{a_{1},a_{2},...,a_{M}\}$, where $a_{j}$ represents the set of cities assigned to the $j$-th agent. Subsequently, a TSP solver $g(a_{j};\mathbf{\mu})$ for each agent $j$ is invoked to order the cities in the allocation $a_{j}$ and return a tour. Finally, the cost of each tour is computed with (2) and the length of the longest tour is the cost of that instance, i.e., $L=\max_{j}D(g(a_{j};\mu))$. The internal process of embedding and the allocation network is below.
@@ IV-C2 Control Variate
$h(d)_{\text{new}}=h(d)+\zeta(c(d)-\mathbb{E}[c(d)]).$ (10)
When constant $\zeta$ is properly chosen, $h(d)_{new}$ has the same expected value but lower variance than $h(d)$, as long as the control variate $c(d)$ is correlated to $h(d)$. We briefly summarize the underlying principle of control variate technique [45] as follows and omit the dependency on $d$ for simplicity.
@@
Let $h$ denotes an RV with an unknown expected value $\mathbb{E}(h)$, and $c$ denotes an RV with a known expected value $\mathbb{E}(c)$. The new RV $h_{new}=h+\zeta(c-\mathbb{E}(c))$ has same expected value but smaller variance, when the constant $\zeta$ is properly chosen and $c$ is correlated with $h$.
@@
$\mathrm{Var}(h_{\text{new}})=\mathrm{Var}(h)+\zeta^{2}\mathrm{Var}(c)+2\zeta\mathrm{Cov}(h,c).$ (11)
Differentiate the above equation w.r.t. $\gamma$ and set the derivative to zero, the optimal choice of $\zeta$ is
$\zeta^{*}=-\frac{\mathrm{Cov}(h,c)}{\mathrm{Var}(c)}.$ (12)
The minimum variance of $h_{\text{new}}$ with $\zeta^{*}$ is $(1-\rho^{2}_{h,c})\mathrm{Var}(h)$ where $\rho$ is the correlation coefficient of $h$ and $c$. Since $\rho\in[-1,1]$, the inequality $\mathrm{Var}(h_{\text{new}})\leq\mathrm{Var}(h)$ always holds with $\zeta=\zeta^{*}$. When $\lVert\rho\rVert\rightarrow 1$, $\mathrm{Var}(h_{\text{new}})\rightarrow 0$. ∎
@@
In practice, $\zeta^{*}$ is usually inaccessible, and thus we cannot arbitrarily choose $c$. Without losing generality, consider the case where $h$ and $c$ are positively correlated. By solving the inequality $\zeta^{2}\mathrm{Var}(c)+2\zeta\mathrm{Cov}(h,c)<0$, we get $\zeta\in(\frac{-2\mathrm{Cov}(h,c)}{\mathrm{Var}(c)},0)$, which indicates that as long as $k$ is set in this range, $\mathrm{Var}(h_{\text{new}})$ is less than $\mathrm{Var}(h)$. When $h$ and $c$ are highly correlated and are of similar ranges, $\mathrm{Cov}(h,c)\approx\mathrm{Var}(c)$, thus $\zeta^{*}\approx-1$, which is our empirical choice of $\zeta$.
Inspired by the control variant gradient estimator (10) we design a lower variance, unbiased gradient estimator using trainable surrogate network $s$. Choose the allocation log-derivative estimator as $h(d)$ in (10), and replace $c(d)$ and $\mathbb{E}[c(d)]$ with the surrogate log-derivative estimator and the actual surrogate network gradient, respectively, the upper-level costs $U_{\theta}$ of our bilevel framework is:
@@
where the constant $\zeta$ in (10) is substituted with $-1$. $L^{\prime}$ is output of the surrogate network (control variate) but is taken as a constant by cutting its gradients. In turn, our gradient for the allocation network parameters $\theta$ is:
@@ V Experiments
We next conduct comprehensive experiments to demonstrate the generalization ability, computational efficiency, and convergence speed of our iMTSP framework. We focus our experiments on large-scale settings ($\geq 400$ cities) because current classic solvers can already efficiently find near-optimal solutions for small-scale problems.
@@ V-A Baselines
To demonstrate the advantages of our method, we compare it with both classic methods and learning-based methods. Specifically, we select a state-of-the-art reinforcement learning (RL)-based model from [14] as a baseline. Their framework can efficiently solve MTSPs with up to a few thousand cities. Using S-batched REINFORCE, it has shown outstanding generalization ability and has a reduced variance for the gradient estimator than the vanilla REINFORCE [23]. The other potential baseline methods [13, 46, 20] either don’t provide source codes or cannot be directly applied to our settings. We also compare iMTSP with a well-known meta-heuristic method implemented by the Google OR-Tools routing library [18, 22, 14]. It provides near-optimal solutions to MTSPs with a few hundred cities and is a popular MTSP baseline [29]. For abbreviation, we will refer to the two methods as an RL baseline and Google OR-Tools. All approaches are tested locally with the same hardware.